OpenStudy (anonymous):

the volume of a cylinder is 980in^3. The radius of the cylinder is 7 in. What is it height : a) 70in b)20in c)140in d)10in

OpenStudy (campbell_st):

well you know \[980 = \pi \times 7^2 \times h\] just solve for h

OpenStudy (campbell_st):

thats interesting... the answer choices don't contain the correct answer...